# Efficient Inference for Large Language Model-based Generative Recommendation

### v2

@@ Abstract
Large Language Model (LLM)-based generative recommendation has achieved notable success, yet its practical deployment is costly particularly due to excessive inference latency caused by autoregressive decoding. For lossless LLM decoding acceleration, Speculative Decoding (SD) has emerged as a promising solution. However, applying SD to generative recommendation presents unique challenges due to the requirement of generating top-$K$ items (i.e., $K$ distinct token sequences) as a recommendation list by beam search. This leads to more stringent verification in SD, where all the top-$K$ sequences from the target LLM must be successfully drafted by the draft model at each decoding step. To alleviate this, we consider 1) boosting top-$K$ sequence alignment between the draft model and the target LLM, and 2) relaxing the verification strategy to reduce trivial LLM calls. To this end, we propose an alignment framework named AtSpeed, which presents the AtSpeed-S optimization objective for top-$K$ alignment under the strict top-$K$ verification. Moreover, we introduce a relaxed sampling verification strategy that allows high-probability non-top-$K$ drafted sequences to be accepted, significantly reducing LLM calls. Correspondingly, we propose AtSpeed-R for top-$K$ alignment under this relaxed sampling verification. Empirical results on two real-world datasets demonstrate that AtSpeed significantly accelerates LLM-based generative recommendation, e.g., near 2$\times$ speedup under strict top-$K$ verification and up to 2.5$\times$ speedup under relaxed sampling verification. The codes and datasets will be released in the near future.
@@ 2 Task Formulation
LLM-based Generative Recommendation. In LLM-based generative recommendation, each item is represented by an item identifier, i.e., a token sequence such as item title (Bao et al., 2023) or learnable token sequence (Zheng et al., 2024), linking the recommendation items to the language space for LLMs to understand user behaviors and recommend items (Lin et al., 2024). Formally, given the user’s historical interactions $\bm{x}$, the well-trained target LLM-based recommender model $\mathcal{M}_{p}$ generates top-$K$ ranked items via beam search, i.e., $\{\bm{y}_{L,i}\}_{i=1}^{K}\leftarrow\mathcal{M}_{p}(\bm{x})$, where $\bm{y}_{L,i}=(y_{1},y_{2},\dots,y_{L})_{i}$ is the item identifier of the $i$-th recommended item of length $L$ [^1] We follow the widely used codebook-based item identifier, due to its promising results and generalization ability on cold-start items (Wang et al., 2024a; Rajput et al., 2023a). The codebook-based identifier will ensure each item has an identifier of the same length.. However, the LLM inference is time-consuming due to the need to perform multiple LLM calls (i.e., forward process) during autoregressive generation. To accelerate the LLM inference, we are motivated to leverage SD for LLM-based generative recommendation for its decoding acceleration without losing accuracy.
@@ 3.1 Alignment for Strict Top-K Verification (AtSpeed-S)
Acceptance Rate. We define the acceptance rate $\beta$ as the probability of the step acceptance, i.e., accepting all the top-$K$ ideal sequences from the draft sequences $\mathcal{Y}_{q}$. With strict top-K verification strategy, for each step, we have $\beta=1$ if ${p}({\bm{y}})\geq{p}(\bm{y}_{K})$ for $\forall\bm{y}\in\mathcal{Y}_{q}$, where $\bm{y}_{K}$ is the sequence that has the K-th highest probability in ${p}$.
@@
where $\mathcal{Y}_{q_{\theta}}=\mathcal{Y}_{1}^{q}\cup\mathcal{Y}_{2}^{q}\dots\cup\mathcal{Y}_{L}^{q}$ aims at maximizing acceptance rate for every beam search step (see detailed derivatives in Appendix A.2). This alignment objective can be further expanded as:
@@
where $c_{t}=(\bm{x},\bm{y}_{<t})$ is the context, $p(y_{K})=p(\bm{y}_{K,t}|\bm{y}_{K,<t})$, and $\mathcal{D}^{\prime}=\{(\bm{x},\mathcal{Y}=\text{Top}K((1-\lambda)\mathcal{M}_{q}(\bm{x})+\lambda\mathcal{M}_{p}))|\bm{x}\sim\mathcal{D}\}$ [^3] We follow (Gu et al., 2024) to adopt the target LLM-mixed sampling for data construction., and $\mathcal{V}$ is the LLM vocabulary.
@@
where $\mathcal{L}_{\text{Rec}}$ is the original loss to train the model for the recommendation tasks and $\alpha$ is the hyper-parameter to control the alignment strength.
@@ 3.2 Alignment for Relaxed Sampling Verification (AtSpeed-R)
Relaxed Sampling Verification. We design the verification strategy for each beam search step $j\in\{1,\dots,\gamma\}$ as follows. Given the drafted results $\mathcal{Y}_{q}\sim q$ [^4] The drafted sequences are sampled from the sequence distribution $q_{j}$, and superscript $j$ is omitted for clarity., for each $\bm{y}\in\mathcal{Y}_{q}$, we accept $\bm{y}$ if ${p}(\bm{y})\geq{q}(\bm{y})$. Otherwise, we reject $\bm{y}$ with probability of $1-\frac{p(\bm{y})}{q(\bm{y})}$ and resample $\bm{y}\sim p^{\prime}=\text{norm}(\max(0,p(\bm{y})-q(\bm{y})))$. Notably, for the non-top-$K$ drafted sequence $\bm{y}$ with a high draft probability (i.e., $q(\bm{y})\geq p(\bm{y})$), this verification strategy ensures they have a certain probability of $\frac{p(\bm{y})}{q(\bm{y})}$ to be accepted. We present more details on the relaxed sampling verification strategy in Algorithm 2 in Appendix. Drawing upon the with-replacement sampling approximation [^5] When the sample size (i.e., the number of recommended items $K$) is much smaller than the total population size (i.e., the vocabulary size), the sampling without replacement can be effectively approximated to the sampling with replacement. See Appendix A.2 for proof., the joint distribution of K output sequences under such verification is approximately equivalent to that of the target LLM as proven in Appendix A.2. Such equivalence ensures that the relaxed sampling verification does not sacrifice much recommendation accuracy (see Section 4.3 for empirical evidence).
@@
The derivation of Eq.(6) can be easily extended from (Leviathan et al., 2023) as in Appendix A.2.
@@
$\theta^{*}:$ $=\mathop{\arg\min}_{\theta\in\Theta}\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}^{\prime}}\sum_{K}\big{[}\frac{1}{|\bm{y}|}\sum_{t=1}^{|\bm{y}|}\text{TVD}(p^{\prime}(\cdot|\bm{x},\bm{y}_{<t}),q^{\prime}(\cdot|\bm{x},\bm{y}_{<t}))\big{]}.$ (7)
@@ 3.3 Inference of AtSpeed
Tree-based Attention. A challenge of AtSpeed inference is the verification inefficiency issue. Using beam search results from every step would largely increase the number of drafted sequences ($\gamma N$ sequences), where some prefix is shared by different beam sequences. As such, self-attention is repeatedly calculated for the same prefix through a single LLM call, thus leading to verification inefficiency, especially with a larger $N$ (e.g., $N=20$). In this work, we leverage the tree-based attention (Miao et al., 2024) to eliminate the repeated calculations via a single flattened sequence, thereby boosting the efficiency of verification. The tree-based attention for the drafted sequences from all steps is illustrated in Figure 2(c).
@@ 4 Experiments
  • RQ1: How does our proposed AtSpeed perform in improving inference efficiency for LLM-based recommendation under both strict top-K and relaxed sampling verification strategies?
  • RQ2: How do different components of AtSpeed affect the decoding acceleration?
  • RQ3: How is the recommendation accuracy of AtSpeed-R under the relaxed sampling verification?
  • RQ4: How do different hyper-parameters affect the performance?
@@ 4.1 Experimental Settings
Datasets and Baselines. To evaluate our proposed framework, we instantiate AtSpeed on a SOTA LLM-based generative recommender model LC-Rec (Zheng et al., 2024) and test on two real-world recommendation datasets from the popular benchmark Amazon review datasets [^6] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/.. 1) Beauty contains user interactions with the beauty products and 2) Games collects the user interactions with the video games. For both datasets, each item has rich textual information such as title and description. Based on the item textual information, we follow LC-Rec to assign each item with an item identifier generated by RQ-VAE (Lee et al., 2022). More details of the datasets can be found in Appendix A.3.
We compare our proposed alignment methods with Supervised Fine-tuning (SFT) and several representative Knowledge Distillation (KD) methods as follows: 1) SFT fine-tunes the draft model with the recommendation loss $\mathcal{L}_{Rec}$ on the recommendation dataset; 2) WordKD (Sanh, 2019; Hinton, 2015) fine-tunes the draft model to align with the token probability of the target LLM by KLD on the recommendation dataset; 3) SeqKD (Kim & Rush, 2016) additionally utilizes the top-K target LLM-generated data to fine-tune the draft model with the recommendation loss $\mathcal{L}_{Rec}$; 4) TVDKD (Wen et al., 2023) aligns the draft model and the target LLM with the symmetric divergence TVD.
Evaluation Metrics. To measure the decoding efficiency, we use walltime speedup@$K$ (WS@$K$) for evaluation. In addition, we report accept step@$K$ (AS@$K$), which is defined as the average number of accepted steps during the SD for each user. As for the recommendation performance, we adopt the widely used metrics Recall@$K$ for evaluation.
Implementation Details. We instantiate AtSpeed on LC-Rec, where the backend LLM is LLaMA-7B (Touvron et al., 2023). For each dataset, we fine-tune the target LLM using the recommendation loss $\mathcal{L}_{\text{Rec}}$ with the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021). As for the draft model, we fully fine-tune a compatible small-sized model LLaMA-68M (Miao et al., 2024) for alignment training, and then utilize the fine-tuned model for SD inference. We set draft length $\gamma=4$, number of recommended items $K=\{1,3,5,10,20\}$, and draft beam size $N=40$. More implementation details are provided in Appendix A.
@@ 4.2 Overall Performance (RQ1)
  • Among the baselines, SeqKD yields competitive results in most cases. This is reasonable since SeqKD 1) aligns the draft model with the target LLM at the sequence level, which fosters the generation of sequences more aligned with the target LLM’s distribution (superior performance than WordKD); and 2) fine-tunes the draft model with target LLM-generated data, which implicitly incorporates the information of valid tokens in item identifiers, thereby alleviating the negative effect from the noisy invalid tokens (superior performance than TVDKD). The competitiveness of SeqKD is also consistent with the observations in NLP tasks (Agarwal et al., 2024).
  • AtSpeed significantly accelerates the LLM-based recommender inference under both strict top-$K$ and relaxed sampling verifications, showing the superiority of our proposed method for strong top-$K$ alignment. In particular, AtSpeed-S under strict top-$K$ verification achieves an average speedup of 1.97$\times$ on Beauty and 1.83$\times$ on Games while obtaining the identical results of the top-$K$ recommended items. Moreover, AtSpeed-R further enhances the efficiency with an average speedup of 2.11$\times$ and 2.05$\times$ on Beauty and Games, respectively. Furthermore, as $K$ increases, the number of accept steps gradually decreases due to the increased difficulty of the $N$-to-$K$ verification. Nevertheless, it is highlighted that the speedup for $K=20$ is still comparable to that for $K=5$ and outperforms the speedup for $K=10$. This is attributed to tree-based attention, which efficiently verifies the drafted sequences (more detailed analysis in Appendix A.3).
Table 1: Overall comparison of walltime speedup (WS@$K$) and accept steps (AS@$K$) between the baselines and AtSpeed instantiated on LC-Rec (LLaMA-7B) across two datasets. For each verification strategy, the best results are highlighted in bold and the second-best results are underlined.
@@
SFT 1.88 1.55 1.43 1.37 1.55 2.19 1.66 1.32 0.66 0.09
WordKD 2.02 1.70 1.58 1.52 1.58 2.33 1.89 1.60 1.03 0.16
TVDKD 1.84 1.55 1.44 1.37 1.57 2.15 1.67 1.31 0.65 0.09
@@
SFT 1.77 1.76 1.80 2.06 2.36 2.15 2.06 2.03 1.99 1.48
WordKD 1.74 1.75 1.81 1.99 2.05 2.11 2.04 2.01 1.87 1.07
TVDKD 1.80 1.77 1.81 2.06 2.35 2.17 2.07 2.03 1.99 1.45
@@
SFT 1.82 1.45 1.43 1.40 1.58 2.13 1.49 1.32 0.91 0.15
WordKD 1.82 1.45 1.31 1.35 1.47 2.13 1.49 1.10 0.80 0.14
TVDKD 1.70 1.33 1.24 1.32 1.50 1.99 1.26 0.95 0.66 0.09

### v3

@@ Abstract
Large Language Model (LLM)-based generative recommendation has achieved notable success, yet its practical deployment is costly particularly due to excessive inference latency caused by autoregressive decoding. For lossless LLM decoding acceleration, Speculative Decoding (SD) has emerged as a promising solution. However, applying SD to generative recommendation presents unique challenges due to the requirement of generating top-$K$ items (i.e., $K$ distinct token sequences) as a recommendation list by beam search. This leads to more stringent verification in SD, where all the top-$K$ sequences from the target LLM must be successfully drafted by the draft model at each decoding step. To alleviate this, we consider 1) boosting top-$K$ sequence alignment between the draft model and the target LLM, and 2) relaxing the verification strategy to reduce trivial LLM calls. To this end, we propose an alignment framework named AtSpeed, which presents the AtSpeed-S optimization objective for top-$K$ alignment under the strict top-$K$ verification. Moreover, we introduce a relaxed sampling verification strategy that allows high-probability non-top-$K$ drafted sequences to be accepted, significantly reducing LLM calls. Correspondingly, we propose AtSpeed-R for top-$K$ alignment under this relaxed sampling verification. Empirical results on two real-world datasets demonstrate that AtSpeed significantly accelerates LLM-based generative recommendation, e.g., near 2$\times$ speedup under strict top-$K$ verification and up to 2.5$\times$ speedup under relaxed sampling verification. The codes and datasets are available at https://github.com/Linxyhaha/AtSpeed.
@@ 2 Task Formulation
LLM-based Generative Recommendation. In LLM-based generative recommendation, each item is represented by an item identifier, i.e., a token sequence such as item title (Bao et al., 2023) or learnable token sequence (Zheng et al., 2024), linking the recommendation items to the language space for LLMs to understand user behaviors and recommend items (Lin et al., 2024a). Formally, given the user’s historical interactions $\bm{x}$, the well-trained target LLM-based recommender model $\mathcal{M}_{p}$ generates top-$K$ ranked items via beam search, i.e., $\{\bm{y}_{L,i}\}_{i=1}^{K}\leftarrow\mathcal{M}_{p}(\bm{x})$, where $\bm{y}_{L,i}=(y_{1},y_{2},\dots,y_{L})_{i}$ is the item identifier of the $i$-th recommended item of length $L$ [^1] We follow the widely used codebook-based item identifier, due to its promising results and generalization ability on cold-start items (Wang et al., 2024a; Rajput et al., 2023a). The codebook-based identifier will ensure each item has an identifier of the same length.. However, the LLM inference is time-consuming due to the need to perform multiple LLM calls (i.e., forward process) during autoregressive generation. To accelerate the LLM inference, we are motivated to leverage SD for LLM-based generative recommendation for its decoding acceleration without losing accuracy.
@@ 3.1 Alignment for Strict Top-K Verification (AtSpeed-S)
Acceptance Rate. We define the acceptance rate $\beta$ as the probability of the step acceptance, i.e., accepting all the top-$K$ ideal sequences from the draft sequences $\mathcal{Y}_{q}$. With strict top-K verification strategy, for each step, we have $\beta=1$ if $\exists\mathcal{Y}^{\prime}_{q}\subseteq\mathcal{Y}_{q}$ such that ${p}({\bm{y}})\geq{p}(\bm{y}_{K})$ for $\forall\bm{y}\in\mathcal{Y}^{\prime}_{q}$ , where $|\mathcal{Y}^{\prime}_{q}|=K$, and $\bm{y}_{K}$ is the sequence that has the K-th highest probability in ${p}$.
@@
where $\mathcal{Y}_{q_{\theta}}=\mathcal{Y}_{1}^{q}\cup\mathcal{Y}_{2}^{q}\dots\cup\mathcal{Y}_{L}^{q}$ aims at maximizing acceptance rate for every beam search step (see detailed derivatives in Appendix A.2.2). This alignment objective can be further expanded as:
@@
where $c_{t}=(\bm{x},\bm{y}_{<t})$ is the context, $p(y_{K})=p(\bm{y}_{K,t}|\bm{y}_{K,<t})$, and $\mathcal{D}^{\prime}=\{(\bm{x},\mathcal{Y}=\text{Top}K((1-\lambda)\mathcal{M}_{q}(\bm{x})+\lambda\mathcal{M}_{p}))|\bm{x}\sim\mathcal{D}\}$ [^3] We follow (Gu et al., 2024) to adopt the target LLM-mixed sampling for data construction to improve training efficiency and generated data quality., and $\mathcal{V}$ is the LLM vocabulary.
@@
where $\mathcal{L}_{\text{Rec}}$ is the original loss to train the model for the recommendation tasks (see Appendix A.3.1) and $\alpha$ is the hyper-parameter to control the alignment strength.
@@ 3.2 Alignment for Relaxed Sampling Verification (AtSpeed-R)
Relaxed Sampling Verification. We design the verification strategy for each beam search step $j\in\{1,\dots,\gamma\}$ as follows. Given the drafted results $\mathcal{Y}_{q}\sim q$ [^4] The drafted sequences are sampled from the sequence distribution $q_{j}$, and superscript $j$ is omitted for clarity., for each $\bm{y}\in\mathcal{Y}_{q}$, we accept $\bm{y}$ if ${p}(\bm{y})\geq{q}(\bm{y})$. Intuitively, if the target LLM is even more confident in generating a high-probability sequence, the candidate sequence is likely to be generated by the LLMs and should be accepted. Otherwise, we reject $\bm{y}$ with probability of $1-\frac{p(\bm{y})}{q(\bm{y})}$ and resample $\bm{y}\sim p^{\prime}=\text{norm}(\max(0,p(\bm{y})-q(\bm{y})))$. Notably, for the non-top-$K$ drafted sequence $\bm{y}$ with a high draft probability (i.e., $q(\bm{y})\geq p(\bm{y})$), this verification strategy ensures they have a certain probability of $\frac{p(\bm{y})}{q(\bm{y})}$ to be accepted. We present more details on the relaxed sampling verification strategy in Algorithm 2 in Appendix. Drawing upon the with-replacement sampling approximation [^5] When the sample size (i.e., the number of recommended items $K$) is much smaller than the total population size (i.e., all sequences for sampling), the sampling without replacement can be effectively approximated to the sampling with replacement. See Appendix A.2.1 for proof., the joint distribution of K output sequences under such verification is approximately equivalent to that of the target LLM as proven in Appendix A.2.1. Such equivalence ensures that the relaxed sampling verification does not sacrifice much recommendation accuracy (see Section 4.2 for empirical evidence).
@@
The derivation of Eq.(6) can be easily extended from (Leviathan et al., 2023) as in Appendix A.2.3.
@@
$\theta^{*}:$ $=\mathop{\arg\min}_{\theta\in\Theta}\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}^{\prime}}K\cdot\big{[}\frac{1}{|\bm{y}|}\sum_{t=1}^{|\bm{y}|}\text{TVD}(p^{\prime}(\cdot|\bm{x},\bm{y}_{<t}),q^{\prime}(\cdot|\bm{x},\bm{y}_{<t}))\big{]}.$ (7)
@@ 3.3 Inference of AtSpeed
Tree-based Attention. A challenge of AtSpeed inference is the verification inefficiency issue. Using beam search results from every step would largely increase the number of drafted sequences ($\gamma N$ sequences), where some prefix is shared by different beam sequences. As such, self-attention is repeatedly calculated for the same prefix through a single LLM call, thus leading to verification inefficiency, especially with a larger $N$ (e.g., $N=20$). In this work, we leverage the tree-based attention (Miao et al., 2024) to eliminate the repeated calculations via a single flattened sequence, thereby boosting the efficiency of verification. The tree-based attention for the drafted sequences from all steps is illustrated in Figure 2(c) and implementation details are presented in Appendix A.3.1.
@@ 4 Experiments
1) RQ1: How does our proposed AtSpeed perform on LLM-based recommendation under strict top-K and relaxed sampling verification strategies? 2) RQ2: How do different components of AtSpeed affect the decoding acceleration? 3) RQ3: How do different hyper-parameters affect the performance?
@@ 4.1 Experimental Settings
Datasets and Baselines. To evaluate our proposed framework, we instantiate AtSpeed on a SOTA LLM-based generative recommender model LC-Rec (Zheng et al., 2024) and test on two real-world recommendation datasets [^6] We also evaluate our methods on MovieLens-1M and Goodreads datasets (refer to Table 6 in Appendix A.3.2). from the popular benchmark Amazon review datasets [^7] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/.. 1) Beauty contains user interactions with the beauty products and 2) Games collects the user interactions with the video games. For both datasets, each item has rich textual information such as title and description. Based on the item textual information, we follow LC-Rec to assign each item with an item identifier generated by RQ-VAE (Lee et al., 2022). More details of the datasets can be found in Appendix A.3.1.
We compare our proposed alignment methods with Supervised Fine-tuning (SFT) and several representative Knowledge Distillation (KD) methods as follows: 1) SFT fine-tunes the draft model with the recommendation loss $\mathcal{L}_{Rec}$ on the recommendation dataset; 2) WordKD (Sanh, 2019; Hinton, 2015) fine-tunes the draft model to align with the token probability of the target LLM by KLD on the recommendation dataset; 3) SeqKD (Kim & Rush, 2016) additionally utilizes the top-K target LLM-generated data to fine-tune the draft model with the recommendation loss $\mathcal{L}_{Rec}$; 4) TVDKD (Wen et al., 2023) aligns the draft model and the target LLM with the symmetric divergence TVD. We also extend a retrieval-based drafting method 5) DARE (Xi et al., 2024) for a comprehensive comparison.
Evaluation Metrics. To measure the decoding efficiency, we use walltime speedup@$K$ (WS@$K$) compared to the standard target LLM inference for evaluation. In addition, we report accept step@$K$ (AS@$K$), which is defined as the average number of accepted steps during the SD for each user. As for the recommendation performance, we adopt the widely used metrics Recall@$K$ and NDCG@$K$ for evaluation.
Implementation Details. We instantiate AtSpeed on LC-Rec, where the backend LLM is LLaMA-7B (Touvron et al., 2023). For each dataset, we fine-tune the target LLM using the recommendation loss $\mathcal{L}_{\text{Rec}}$ with the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021). As for the draft model, we fully fine-tune a compatible small-sized model LLaMA-68M (Miao et al., 2024) for alignment training, and then utilize the fine-tuned model for SD inference. We set draft length $\gamma=4$, number of recommended items $K=\{1,3,5,10,20\}$, and draft beam size $N=40$. More implementation details are provided in Appendix A.3.1.
@@ 4.2 Overall Performance (RQ1)
  • Among the baselines, SeqKD yields competitive results in most cases. This is reasonable since SeqKD 1) aligns the draft model with the target LLM at the sequence level, which fosters the generation of sequences more aligned with the target LLM’s distribution (superior performance than WordKD); and 2) fine-tunes the draft model with target LLM-generated data, which implicitly incorporates the information of valid tokens in item identifiers, thereby alleviating the negative effect from the noisy invalid tokens (superior performance than TVDKD). The competitiveness of SeqKD is also consistent with the observations in NLP tasks (Agarwal et al., 2024). The less satisfying performance of DARE is reasonable since the uniformly retrieved valid sequences might not be aligned with the target LLM specifically on top-$K$ sequences.
  • AtSpeed significantly accelerates the LLM-based recommender inference under both strict top-$K$ and relaxed sampling verifications, showing the superiority of our proposed method for strong top-$K$ alignment. In particular, AtSpeed-S under strict top-$K$ verification achieves an average speedup of 1.97$\times$ on Beauty and 1.83$\times$ on Games while obtaining the identical results of the top-$K$ recommended items. Moreover, AtSpeed-R further enhances the efficiency with an average speedup of 2.11$\times$ and 2.05$\times$ on Beauty and Games, respectively. Furthermore, as $K$ increases, the number of accept steps gradually decreases due to the increased difficulty of the $N$-to-$K$ verification. Nevertheless, it is highlighted that the speedup for $K=20$ is still comparable to that for $K=5$ and outperforms the speedup for $K=10$. This is attributed to tree-based attention, which efficiently verifies the drafted sequences (more detailed analysis in Appendix A.3.2).
  • The ranking performance under strict top-$K$ verification is lossless because we only accept the perfectly matched top-$K$ beam sequences to obtain identical recommendation results. For relaxed sampling verification, the ranking performance across different alignment methods only shows limited performance drops compared to the target LLM’s top-$K$ results, i.e., performance under strict top-$K$ verification. This meets our expectations since the sampling-based verification ensures the approximately equivalent distribution between the SD output and target LLM output (refer to Appendix A.3.2 for additional accuracy results and analysis).
Table 1: Overall comparison of walltime speedup (WS@$K$) and accept steps (AS@$K$) between the baselines and AtSpeed instantiated on LC-Rec (LLaMA-7B) across two datasets. “Avg” is calculated across $K$ from $1$ to $20$ (full results are presented in Table 4 in Appendix A.3.2). For each verification strategy, the best results are highlighted in bold and the second-best results are underlined.
@@
Verification Method WS@5 WS@10 WS@20 Avg WS AS@5 AS@10 AS@20 Avg AS Recall@5 NDCG@5
Strict Top-$\bm{K}$ DARE 1.06 1.15 1.48 1.18 0.26 0.05 0.00 0.31 0.0056 0.0051
SFT 1.43 1.37 1.55 1.56 1.32 0.66 0.09 1.18 0.0056 0.0051
WordKD 1.58 1.52 1.58 1.68 1.60 1.03 0.16 1.40 0.0056 0.0051
TVDKD 1.44 1.37 1.57 1.55 1.31 0.65 0.09 1.17 0.0056 0.0051
@@
Relaxed Sampling DARE 1.70 1.53 1.69 1.73 1.97 1.14 1.00 1.62 0.0059 ↑0.0003 0.0044 ↓0.0007
SFT 1.80 2.06 2.36 1.95 2.03 1.99 1.48 1.94 0.0057 ↑0.0001 0.0041 ↓0.0010
WordKD 1.81 1.99 2.05 1.87 2.01 1.87 1.07 1.82 0.0066 ↑0.0010 0.0043 ↓0.0008
TVDKD 1.81 2.06 2.35 1.96 2.03 1.99 1.45 1.94 0.0057 ↑0.0001 0.0045 ↓0.0006
@@
Verification Method WS@5 WS@10 WS@20 Avg WS AS@5 AS@10 AS@20 Avg AS Recall@5 NDCG@5
Strict Top-$\bm{K}$ DARE 0.99 1.13 1.44 1.09 0.00 0.00 0.00 0.03 0.0074 0.0065
SFT 1.43 1.40 1.58 1.53 1.32 0.91 0.15 1.20 0.0074 0.0065
WordKD 1.31 1.35 1.47 1.48 1.10 0.80 0.14 1.13 0.0074 0.0065
TVDKD 1.24 1.32 1.50 1.42 0.95 0.66 0.09 0.99 0.0074 0.0065
SeqKD 1.60 1.46 1.77 1.71 1.67 1.05 0.76 1.55 0.0074 0.0065
@@
Relaxed Sampling DARE 1.68 1.19 1.42 1.51 1.96 0.37 0.00 1.27 0.0076 ↑0.0002 0.0065 ↑0.0000
SFT 1.84 1.97 1.69 1.86 2.05 1.89 0.58 1.78 0.0073 ↓0.0001 0.0060 ↓0.0005
WordKD 1.78 1.84 1.56 1.76 1.99 1.68 0.25 1.63 0.0072 ↓0.0002 0.0058 ↓0.0007
TVDKD 1.81 1.90 1.55 1.80 2.02 1.80 0.29 1.69 0.0069 ↓0.0005 0.0061 ↓0.0004
SeqKD 1.90 2.03 2.05 1.95 2.10 1.98 1.22 1.93 0.0071 ↓0.0003 0.0059 ↓0.0006